# Iterative Utility Judgment Framework via LLMs Inspired by Relevance in Philosophy

### v1

@@ 4.2 Evaluation metrics
For the utility judgments task, we evaluate the results of judgments using Precision, Recall, and F1. For the ranking task, we use the normalized discounted cumulative gain (NDCG) Järvelin and Kekäläinen (2017) to evaluate the ranking performance. For the answer generation task, we use the standard exact match metric (EM) and F1, as used by Rajpurkar et al. (2016).
@@ 4.3 LLMs
we conduct our experiments using several representative LLMs, i.e., (i) the well-known ChatGPT OpenAI (2022) (we use the gpt-3.5-turbo-1106 version), (ii) Mistral Jiang et al. (2023a) (the Mistral-7B-Instruct-v0.2 version after instruction tuning), and (iii) Llama 3 Meta (2024) (the Meta-Llama-3-8B-Instruct version promoted with instruction tuning based on the vanilla Llama 3). To ensure the reproducibility of the experiments, the temperature for all experiments is set to 0.
4.4 Retrievers
We use two representative retrievers to gather candidate passages in for utility judgments. Following with previous works Zhang et al. (2024); Sun et al. (2023), we employ RocketQAv2 Ren et al. (2021) and BM25 Robertson et al. (2009) for the NQ dataset and retrieval datasets(i.e., TREC and WebAP datasets), respectively. Based on the retrieval results to build the we have two settings: (i) For TREC and WebAP datasets, we select the top-20 BM25 retrieval results. If these do not include passages with utility, we replaced the last one with a utility-annotated passage. (ii) For the NQ dataset, we use the top-10 dense retrieval results to form the candidate list , following the GTU setting of Zhang et al. (2024).
@@ 4.5 Baselines
We choose the following baselines for comparison:
Sing-shot utility judgments: (i) Vanilla: Ask LLMs to provide utility judgments based on the instruction directly. (ii) UJ-ExpA: Utility judgments facilitated with explicit answers, which is shown to be effective in Zhang et al. (2024). (iii) UJ-ImpA: We utilize the necessary information that is necessary for answering the question to help with the utility judgments task. More details can be found in Section 3.2.
-sampling is proposed by Zhang et al. (2024): In addition to the original input order, -sampling method randomizes the order of input passage list times and aggregates the utility judgement results through voting. The output of the -sampling method each time contains explicit answers and utility judgments. Our iterative framework outputs both answers and utility judgments at each iteration. To ensure fairness in calculating the parameters of LLM, we set . The detailed comparison of the parameter calculations of the two is in Appendix G. The -sampling method requires outputs of LLMs, which can be more expensive for practical applications involving the LLMs. Moreover, it is difficult to determine the stability of the utility judgments due to the sampling technique.
| Method | TREC | WebAP | ||||||||||
| Listwise | Pointwise | Listwise | Pointwise | |||||||||
| Mis. | Lla. | Cha. | Mis. | Lla. | Cha. | Mis. | Lla. | Cha. | Mis. | Lla. | Cha. | |
| Vanilla | 45.67 | 49.39 | 55.19 | 45.11 | 47.64 | 49.84 | 20.79 | 21.79 | 28.43 | 23.05 | 25.09 | 26.85 |
| UJ-ExpA | 54.10 | 52.83 | 57.49 | 43.53 | 53.73 | 48.09 | 27.94 | 26.99 | 30.50 | 25.27 | 29.25 | 27.44 |
| UJ-ImpA | 48.29 | 48.22 | 56.18 | 48.31 | 50.20 | 48.83 | 25.06 | 26.22 | 29.89 | 28.35 | 25.29 | 26.32 |
| -sampling | 52.31 | 52.68 | 60.49 | - | - | - | 30.16 | 28.97 | 31.49 | - | - | - |
| ITEM-As ExpA (1) | 53.78 | 53.66 | 62.52 | 49.44 | 52.09 | 53.61 | 29.76 | 27.50 | 36.89 | 29.10 | 31.08 | 32.02 |
| ITEM-As ExpA (3) | 54.86 | 56.03 | 63.18 | 51.74 | 52.46 | 55.74 | 31.65 | 29.32 | 39.57 | 30.50 | 32.67 | 31.43 |
| ITEM-As ImpA (1) | 49.39 | 53.73 | 58.11 | 46.01 | 53.68 | 54.61 | 26.06 | 25.59 | 34.97 | 28.28 | 30.53 | 29.34 |
| ITEM-As ImpA (3) | 52.05 | 55.14 | 60.56 | 46.59 | 53.76 | 54.90 | 28.36 | 26.10 | 40.78 | 30.13 | 29.64 | 32.54 |
| ITEM-ARs (1) | 52.34 | 48.97 | 62.00 | - | - | - | 35.50 | 31.44 | 36.58 | - | - | - |
| ITEM-ARs (3) | 56.27 | 52.10 | 61.37 | - | - | - | 37.06 | 29.08 | 38.58 | - | - | - |
@@ 5 Experimental Results of LLMs
Our framework consists of three tasks: utility judgments, ranking of topical relevance, and answer generation. This section will present the performance of each task within our framework. By default, the pseudo answer is the explicit answer in all experiments, if not specified otherwise.
5.1 Utility Judgments on Retrieval Datasets
Tabel 1 shows the F1 performance on the TREC and WebAP datasets using three LLMs. Since ITEM-Ar has poor F1 performance in utility judgments (refer to Table 12 in Appendix B.1 for details), we restrict our experiments to ITEM-As and ITEM-ARs in this section.
@@ 5.1 Utility Judgments on Retrieval Datasets
| T | W | T | W | T | W | |
| BM25 | 58.69 | 21.89 | 58.69 | 21.89 | 58.69 | 21.89 |
@@
| ITEM-Ar (1) | 70.57 | 37.11 | 73.95 | 40.89 | 80.79 | 50.30 |
| ITEM-Ar (3) | 74.27 | 43.80 | 77.34 | 45.88 | 83.12 | 51.61 |
ITEM with a Single Iteration vs. Baselines. All three LLMs using our ITEM with a single iteration generally outperform the single-shot utility judgments on both datasets and may even surpass the -sampling method. For example, ChatGPT on the TREC dataset using our ITEM-As ExpA and ImpA in the listwise approach improve the F1 performance by 8.7% and 3.4% over UJ-ExpA and UJ-ImpA, respectively. On the WebAP dataset, Mistral and Llama 3 use single-iteration ITEM-ARs to outperform all baselines and ITEM-As. This might be due to that better topical relevance ranking leads to better utility judgments, which is also better than -sampling in terms of efficacy and efficiency.
ITEM with Multiple Iterations vs. ITEM with Single Iteration. All LLMs using our ITEM-A and ITEM-AR generally demonstrate improved performance with multiple iterations compared to single iterations. For instance, on the WebAP dataset, Mistral, Llama 3, and ChatGPT (using our ITEM-A ExpA) improved their F1 scores in the listwise approach by 6.4%, 6.6%, and 7.3%, respectively, after multiple iterations. However, in some specific cases, multiple iterations may not outperform single iterations, likely due to the unpredictable nature of zero-shot settings. Moreover, our method achieves state-of-the-art performance compared to all baselines by leveraging the iterative framework. The performance improvement from multiple iterations underscores the significance of iterative interaction and further supports Schutz’s interactive framework.
ExpA vs. ImpA. For the single-shot baselines, explicit answers generally outperform implicit answers in utility judgments when using the listwise approach, while the opposite is true for the pointwise approach. This may be because pointwise inputs provide less information, making explicit answers less effective, whereas implicit answers better address the question’s information needs. Unlike the pseudo answer generated by the input passage of the single-shot pointwise approach, our ITEM’s pointwise approach generates the pseudo answer based on the results of utility judgments from the previous round. Therefore, after multiple iterations, ITEM shows greater improvement with explicit answers compared to implicit answers in both input approaches in most cases.
Comparisons between LLMs. ChatGPT outperforms other LLMs on both datasets using both input approaches. For instance, on the TREC dataset, ChatGPT achieves an F1 improvement of 15.2% and 12.8% using the listwise approach after multiple iterations of ITEM-As ExpA method, compared to Mistral and Llama 3, respectively.
Listwise vs. Pointwise. The average performance of utility judgments for LLMs is better with the listwise approach than with the pointwise approach. For example, Mistral shows a 6.0% F1 improvement with the listwise approach over the pointwise approach after multiple iterations of the ITEM-As ExpA method on the TREC dataset. More detailed performance metrics, i.e., Precision, Recall, F1, are provided in Appendix B.1.
5.2 Ranking Performance on Retrieval Datasets
@@
We also assess whether the ranking of topical relevance has been improved within ITEM. In terms of topical relevance, the utility ranking results produced by ITEM-Ar and the relevance ranking results yielded from ITEM-ARs. From Table 2, we can observe that: (i) Our ITEM with single iteration significantly improves the ranking of topical relevance performance compared to the RankGPT. For instance, relevance ranking outperforms RankGPT in NDCG@5 by 2.1% on the TREC dataset and 27.7% on the WebAP dataset using the Mistral. The performance improvement may stem from the interaction between tasks. (ii) After iterations, relevance ranking, and utility ranking performances have been improved on both datasets and all LLMs. The topical relevance ranking benefits from our dynamic iterative framework, confirming Schutz’s theory of dynamic iterative interaction. (iii) Experiments show that utility ranking generally outperforms relevance ranking, which confirms the effectiveness of utility in the ranking of topical relevance.
@@ 5.3 Results of Answer Generation
From Table 1, we find that the listwise approach generally outperforms the pointwise approach for utility judgments. Consequently, our NQ dataset experiments utilize only the listwise approach.
@@
The following observations can be made from Table 3: (i) The -sampling method has the best performing, especial on F1 score, for all the baselines. (ii) ITEM outperforms baselines on all metrics on all LLMs (except for the EM score of Llama 3), indicating that ITEM can help the LLMs to find better evidence for generating answers. (iii) The results from Llama 3 have better answer performance than ChatGPT, and we suspect that Llama 3 has been over-tuned on the NQ dataset.
@@ 6 Further Analyses
6.1 Effect of Iteration Numbers
Figure 6 illustrates the performance of the LLMs in making utility judgments under ITEM-As with varying maximum iteration rounds on the TREC dataset. Figure 7 illustrates the ranking of topical relevance performance in terms of values. Since the NQ dataset lacks annotations for passage utility, we automatically annotated the retrieved passages based on the utility standard, as detailed in Appendix D. Table 4 displays the results for utility judgments and answer generation using Mistral under ITEM-As on the NQ dataset.
| ITEM-As | ITEM-ARs | |||||
| Utility | Answer | Utility | Answer | |||
| F1 | EM | F1 | F1 | EM | F1 | |
| =1 | 28.83 | 32.98 | 49.00 | 24.96 | 33.30 | 49.26 |
| =2 | 29.02 | 34.31 | 50.08 | 25.87 | 33.57 | 49.16 |
| =3 | 28.90 | 33.73 | 49.63 | 26.17 | 33.40 | 49.27 |
| =4 | 29.01 | 34.21 | 50.07 | 26.54 | 33.46 | 49.24 |
| =5 | 28.91 | 33.78 | 49.63 | 26.48 | 33.89 | 49.58 |
As can be seen from Figure 6, Figure 7 and Table 4, varying the value of affects the performance of utility judgments, ranking of topical relevance and answer generation. However, the performance when is consistently better than when . Moreover, utility ranking generally outperforms relevance ranking, which confirms the effectiveness of utility in the ranking of topical relevance. Further details on the experiments with different values are provided in Appendix B.1.
In addition, we evaluate two ranking performances of ITEM-ARr, with the experimental results shown in Table 5. Under the ITEM-ARr framework, relevance ranking is followed by utility ranking, resulting in further improved utility ranking performance compared to relevance ranking. However, as seen in Table 2 and Table 5, performing ranking twice in the same iteration may not yield better ranking results than performing utility ranking once in the iteration.
| NDCG@5 | NDCG@10 | NDCG@20 | Utility-F1 | |
| 1 | 71.29 / 72.77 | 72.90 / 74.96 | 84.56 / 85.75 | 43.13 |
| 2 | 72.54 / 70.99 | 74.81 / 73.76 | 85.77 / 85.28 | 40.21 |
| 3 | 72.07 / 74.14 | 74.14 / 76.63 | 85.53 / 86.57 | 45.67 |
| 4 | 71.02 / 71.06 | 74.30 / 74.03 | 85.09 / 85.16 | 43.82 |
| 5 | 72.26 / 70.12 | 75.83 / 72.59 | 85.88 / 84.77 | 44.10 |
6.2 Iteration Stop Conditions
The stopping condition takes into account the answer performance in our ITEM in addition to the utility judgments. we compute the ROUGE-L Lin (2004) score of the answer in two iterations, and stop the iteration if the ROUGE-L of and is greater than . The utility judgments performance of different iteration stop conditions are shown in Figure 8. More details of experiments on iteration stop conditions are shown in Appendix B.2.
The results show that using different stopping conditions does affect the performance of utility judgments. However, using the answer as a stopping condition, different models on different datasets may need to look for different , which is not very flexible.
@@ 7 Case Study
Table 16 and Table 17 in Appendix C present two good cases (demonstrating improved utility judgments after multiple iterations) from the TREC dataset using Mistral on ITEM-As. In Table 16, the first pseudo answer, though relatively correct, includes irrelevant information, leading to a misjudgment of “Passage-2” as “utility”. “Passage-2” does not address the question of what tissue bronchioles belong to but instead discusses the type of muscles that control the size of the bronchioles in the lungs. Based on the results of the first round of utility judgments, the second round of the pseudo answer is more accurate and free of irrelevant content. Consequently, all three passages in the second round of utility judgments have utility in answering the question. In Table 17, the first pseudo answer is correct, but two irrelevant passages are judged as “utility”. The second pseudo answer, with slight rewording, results in all passages being correctly identified as “utility”.

### v2

@@ 4.2 Evaluation metrics
For the utility judgments task, we evaluate the results of judgments using Precision, Recall, and micro F1. For the ranking task, we use the normalized discounted cumulative gain (NDCG) Järvelin and Kekäläinen (2017) to evaluate the ranking performance. For the answer generation task, we use the standard exact match (EM) metric and F1.
@@ 4.3 LLMs
We conduct our experiments using several representative LLMs, i.e., (i) ChatGPT OpenAI (2022) (we use the gpt-3.5-turbo-1106 version), (ii) Mistral Jiang et al. (2023a) (the Mistral-7B-Instruct-v0.2 version), and (iii) Llama 3 Meta (2024) (the Meta-Llama-3-8B-Instruct version). To ensure the reproducibility of the experiments, the temperature for all experiments is set to 0.
| Method | WebAP | TREC DL | ||||||||||
| Listwise | Pointwise | Listwise | Pointwise | |||||||||
| M | L | C | M | L | C | M | L | C | M | L | C | |
| Vanilla | 20.79 | 21.79 | 28.43 | 23.05 | 25.09 | 26.85 | 45.67 | 49.39 | 55.19 | 45.11 | 47.64 | 49.84 |
| UJ-ExpA | 27.94 | 26.99 | 30.50 | 25.27 | 29.25 | 27.44 | 54.10 | 52.83 | 57.49 | 43.53 | 53.73 | 48.09 |
| UJ-ImpA | 25.06 | 26.22 | 29.89 | 28.35 | 25.29 | 26.32 | 48.29 | 48.22 | 56.18 | 48.31 | 50.20 | 48.83 |
| -sampling | 30.16 | 28.97 | 31.49 | - | - | - | 52.31 | 52.68 | 60.49 | - | - | - |
| ITEM-As ExpA (1) | 29.76 | 27.50 | 36.89 | 29.10 | 31.08 | 32.02 | 53.78 | 53.66 | 62.52 | 49.44 | 52.09 | 53.61 |
| ITEM-As ImpA (1) | 26.06 | 25.59 | 34.97 | 28.28 | 30.53 | 29.34 | 49.39 | 53.73 | 58.11 | 46.01 | 53.68 | 54.61 |
| ITEM-ARs ExpA(1) | 35.50 | 31.44 | 36.58 | - | - | - | 52.34 | 48.97 | 62.00 | - | - | - |
| ITEM-As ExpA (3) | 31.65 | 29.32 | 39.57 | 30.50 | 32.67 | 31.43 | 54.86 | 56.03 | 63.18 | 51.74 | 52.46 | 55.74 |
| ITEM-As ImpA (3) | 28.36 | 26.10 | 40.78 | 30.13 | 29.64 | 32.54 | 52.05 | 55.14 | 60.56 | 46.59 | 53.76 | 54.90 |
| ITEM-ARs ExpA(3) | 37.06 | 29.08 | 38.58 | - | - | - | 56.27 | 52.10 | 61.37 | - | - | - |
@@ 4.4 Baselines
We utilize the following baselines on the utility judgments task and question answering performance based on the utility judgment results:
Single-shot utility judgments. (i) Vanilla: Ask LLMs to provide utility judgments based on the instruction directly. (ii) UJ-ExpA: Utility judgments and provide explicit answers simultaneously through a single output, which is shown to be effective in Zhang et al. (2024). (iii) UJ-ImpA: Utility judgments and provide implicit answers that are necessary to answer the question through a single output.
-sampling. Zhang et al. (2024) proposed -sampling to alleviate the sensitivity of LLMs to input order. Specifically, the -sampling method randomizes the order of the input passage list times in addition to the original input and aggregates the utility judgment results through voting. For fair comparison, we use the , more details are on Appendix E.
To evaluate the effectiveness of the proposed ITEM framework in ranking tasks, we are using a verberlized ranking. Therefore, we also employ another verberlized ranking method, i.e., RankGPT Sun et al. (2023) as our main baseline, which uses the LLMs to directly rank input passages based on their relevance to the query.
| Method | Llama3 8B | ChatGPT | ||
| Listwise | Pointwise | Listwise | Pointwise | |
| Vanilla | 43.38 | 28.55 | 59.37 | 35.31 |
| UJ-ExapA | 47.07 | 39.32 | 66.13 | 37.17 |
| UJ-ImpA | 43.31 | 38.72 | 57.40 | 37.29 |
| k-sampling | 49.20 | - | 71.17 | - |
| ITEM-As-ExpA(1) | 49.26 | 47.52 | 72.44 | 54.89 |
| ITEM-As-Imp(1) | 47.47 | 37.98 | 68.92 | 43.17 |
| ITEM-ARs-ExpA(1) | 50.77 | - | 74.43 | - |
| ITEM-As-ExpA(3) | 49.73 | 48.90 | 73.55 | 55.45 |
| ITEM-As-Imp(3) | 48.03 | 38.34 | 69.68 | 43.58 |
| ITEM-ARs-ExpA(3) | 51.22 | - | 76.34 | - |
@@ 5 Experimental Results of LLMs
This section will present the performance of each task within our ITEM framework. By default, the pseudo answer is the explicit answer in all experiments, if not specified otherwise.
5.1 Utility Judgments Results
Table 1 shows the micro F1 performance on the TREC DL and WebAP datasets using three LLMs. Further, we utilize a better-performing open-source LLM, i.e., Llama-3.1-8B, and a closed LLM, i.e., ChatGPT, to conduct experiments on GTI-NQ, as shown in Table 2. Since ITEM-Ar and ITEM-ARr have poor F1 performance in utility judgments (refer to Table 12 in Appendix A.4 for details), we restrict our experiments to ITEM-As and ITEM-ARs in this section.
ITEM with a Single Iteration vs. Baselines. All LLMs using our ITEM with a single iteration generally outperform the single-shot utility judgments on three datasets and may even surpass the -sampling method. For example, ChatGPT on the TREC DL dataset using our ITEM-As ExpA and ImpA in the listwise approach improve the F1 performance by 8.7% and 3.4% over UJ-ExpA and UJ-ImpA, respectively. Explicit generation of pseudo-answers by LLMs enhances their performance in utility judgment tasks, highlighting the importance of task interaction. Moreover, concurrent execution of answer generation and utility judgment within a single inference cycle yields inferior performance compared to sequential task execution through separate reasoning phases.
ITEM with Multiple Iterations vs. ITEM with Single Iteration. All LLMs using our ITEM-A and ITEM-AR generally demonstrate improved performance with multiple iterations compared to single iterations on all three datasets. For instance, on the WebAP dataset, Mistral, Llama 3, and ChatGPT (using our ITEM-A ExpA) improved their F1 scores in the listwise approach by 6.4%, 6.6%, and 7.3%, respectively, after multiple iterations. Moreover, our method achieves state-of-the-art performance compared to all baselines by leveraging the iterative framework. The performance improvement from multiple iterations underscores the significance of iterative interaction and further supports Schutz’s interactive framework. However, in some specific cases, multiple iterations may not outperform single iterations, likely due to the unpredictable nature of zero-shot settings. ChatGPT outperforms other LLMs on all datasets using both input approaches.
ITEM-As vs. ITEM-ARs. In our utility-emphasized iterative RAG framework, ITEM-As and ITEM-ARs are the two major methods we propose. From Table 1, we find that ITEM-ARs works better than ITEM-As most of the time for complex questions (WebAP, all the questions are non-factoid) and the complex candidate passage list (GTI-NQ, containing different types of passage), indicating complicated question or passage list need more components in the loop. For TREC DL, which contains factoid questions, we find that ITEM-ARs is worse than ITEM-As most times on TREC DL. This is reasonable since factoid questions are relatively easier to answer and may not need more components involved in the iteration.
Listwise vs. Pointwise. The general performance of utility judgments for LLMs is better with the listwise approach than with the pointwise approach. The primary rationale lies in the listwise approach exposes the LLM to broader contextual information, thereby facilitating more effective interaction during the LLMs in judging the passages’ utility.
@@ 5.1 Utility Judgments Results
| TREC | WebAP | TREC | WebAP | TREC | WebAP | |
| 58.69 | 21.89 | 58.69 | 21.89 | 58.69 | 21.89 | |
@@
| Method | Llama 3 | ChatGPT | ||
| @5 | @10 | @5 | @10 | |
| 29.46 | 45.26 | 29.46 | 45.26 | |
| RankGPT | 71.50 | 74.05 | 77.27 | 78.64 |
| ITEM-Ar(m=1) | 74.36 | 76.91 | 85.99 | 87.26 |
| ITEM-ARs(m=1) | 75.46 | 77.75 | 84.54 | 85.14 |
| ITEM-Ar(m=3) | 75.95 | 78.18 | 87.48 | 88.47 |
| ITEM-ARs(m=3) | 76.38 | 78.56 | 85.95 | 86.39 |
5.2 Ranking Performance
@@
We also assess whether the ranking performance has been improved within ITEM on retrieval datasets (Table 3) and utility judgments benchmark (Table 4). In terms of ranking performance, we consider two rankings: relevance ranking (ITEM-ARs) and utility ranking (ITEM-Ar). We can observe that: (i) Our ITEM with single iteration significantly improves the ranking of topical relevance performance compared to the RankGPT. For instance, relevance ranking outperforms RankGPT in NDCG@5 by 2.1% on the TREC dataset. The performance improvement may stem from the interaction between tasks. (ii) After iterations, relevance and utility ranking performance have been improved on all datasets and all LLMs. The ranking benefits from our dynamic iterative framework, confirming Schutz’s theory of dynamic iterative interaction. (iii) From Table 3&4, we can find that ITEM-ARs is generally better than ITEM-Ar when . However, when , it may have the opposite performance. The possible reason is that when is small, the answer is very good, and utility is more dependent on the answer than relevance. However, as iterations occur, the quality of the answer is better, and utility performance is gradually improved, while relevance does not have as obvious improvement effect as utility in iterations. This indicates that reranking based on utility performs better than including more components in the loop, but still using relevance as the ranking criterion. These findings further confirm the importance of the concept of utility in RAG.
@@ 5.3 Results of Answer Generation
In the answer generation task, the results of utility judgments are fed to LLMs for answer generation. We use the factoid QA dataset (i.e., NQ) for answer generation evaluation, as shown in Table 5. From Table 1&2, we find that the listwise approach generally outperforms the pointwise approach for utility judgments. Consequently, our answer generation experiments utilize only the listwise utility judgments.
@@
The following observations can be made from Table 5: (i) ITEM outperforms baselines on all metrics on all LLMs (except for the EM score of Llama 3), indicating that ITEM can help the LLMs to find better evidence for generating answers. (ii) Similar to Table 1&2, when the , ITEM-ARs performs better than ITEM-As, which shows the importance of relevance reranking in ITEM. However, as the number of iterations increases, ITEM-As performs better than ITEM-ARs. We are keen to discern the optimal choice for different tasks: 1) More components and more iterations are not always needed, especially for simpler tasks; 2) Fewer iterations with numerous components, or increased iterations with few components.
@@ 6 Further Analyses
Iteration Rounds. Figure 7 shows the performance of (a): utility judgments under ITEM-As and (b): ranking with varying maximum iteration rounds .
We observe the following: 1) Varying the value of affects the performance of utility judgments, ranking. 2) Based on empirical observations balancing the cost and performance, the was operationally configured with distinct values for different question types on utility judgments (=3 in our paper on all experiments for fair comparison): =2 for factoid questions, whereas =3 was implemented for non-factoid questions in practical applications. 3) Utility ranking generally outperforms relevance ranking, which confirms the effectiveness of utility in the ranking task.
Iteration Stop Conditions. In addition to utility judgments, we also consider the pseudo answer generation performance in ITEM as a stopping condition. Specifically, we calculate the ROUGE-L score Lin (2004) of the answer in two iterations and stop the iteration if the ROUGE-L of and is greater than . The utility judgments performance of different iteration stop conditions are shown in Figure 8. The results show that using different stopping conditions affects the performance of utility judgments. However, using the answer as a stopping condition, different LLMs on different datasets may need to look for different , which is not very flexible.
7 Inference Efficiency
Table 14 in Appendix G shows the inference efficiency analysis of our ITEM. Our iteration framework surpasses -sampling in computational efficiency during inference. The proposed approach demonstrates potential for large-scale retrieval data annotation, where ITEM-As emerges as an optimal solution by simultaneously enhancing both performance and operational efficiency, thereby facilitating utility annotation for retrieval systems.
@@ 8 Case Study
Figure 9 in Appendix B presents two cases from the TREC DL dataset using Mistral under ITEM-As. For the first question in Figure 9, the first pseudo-answer, though relatively correct, includes irrelevant information, leading to a misjudgment of “Passage-2” as “utility”. Based on the results of the first round of utility judgments, the second round of the pseudo-answer is more accurate and free of irrelevant content. Consequently, all three passages in the second round of utility judgments have utility in answering the question. For the second question in Table 9, the first pseudo-answer is correct, but two passages without utility are judged as “utility”. The second pseudo-answer, with slight rewording, results in all selected passages being correct.